OpenStudy (anonymous):

I cannot figure out how to eliminate this parameter! x = t^2 and y = 2ln(t), eliminate the parameter and write the corresponding rectangular equation.

OpenStudy (loser66):

solve from y =lnt^2=lnx

OpenStudy (loser66):

that's it

OpenStudy (anonymous):

Uhh... Could you explain it?

OpenStudy (loser66):

x= t^2 y =2lnt = ln t^2 replace t^2 =x so y= ln x.

OpenStudy (anonymous):

Oh! I see now.

OpenStudy (loser66):

you don't know that 2 ln t = ln (t)^2?

OpenStudy (anonymous):

I forgot... Lol. I can;t figure out that problem and one other one.

OpenStudy (loser66):

ok

OpenStudy (anonymous):

Thanks. :D

OpenStudy (loser66):

w